What is the hydroxide-ion concentration of an aqueous solution having a hydronium-ion concentration of 1.905 × 10⁻⁵ mol/L?
1
Expert's answer
2020-05-15T10:05:21-0400

Solution:

[H3O+] = 1.905×10−5 mol/L

The hydroxide ion concentration can be calculated using the Kw constant:

Kw = [H3O+]×[OH] = 1.0×10-14

Therefore,

[OH] = Kw / [H3O+]

[OH] = (1.0×10−14) / (1.905×10−5) = 5.249×10−10

[OH] = 5.249×10−10 M


Answer: [OH] = 5.249×10−10 M